OpenStudy (inhaleof):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31? x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left. A number line with a closed circle on -9 and shading to the right.

OpenStudy (amoodarya):

8x - (5x + 4) greater than or equal to -31 8x - (5x + 4) >= -31

OpenStudy (amoodarya):

8x - (5x + 4) >= -31 8x-5x-4>=-31 3x>=4-31 3x>=-27 /3 x>-9

OpenStudy (texaschic101):

amodayra solved the inequality for you....except for the = sign. It should be : x >= -9 (basically this is saying x is greater then or equal to -9) Here is a little tip ...if there is an equal sign, then there is a closed circle. And if it is > (greater then), shading goes to the right.